# 22affiliationtext: Booth School of Business

### v1

@@ Abstract
In designing generative models, it is commonly believed that in order to learn useful latent structure, we face a fundamental tension between expressivity and structure. In this paper we challenge this view by proposing a new approach to training arbitrarily expressive generative models that simultaneously learn disentangled latent structure. This is accomplished by adding a simple decoder-only module to the head of an existing decoder block that can be arbitrarily complex. The module learns to process concept information by implicitly inverting linear representations from an encoder. Inspired by the notion of intervention in causal graphical models, our module selectively modifies its architecture during training, allowing it to learn a compact joint model over different contexts. We show how adding this module leads to disentangled representations that can be composed for out-of-distribution generation. To further validate our proposed approach, we prove a new identifiability result that extends existing work on identifying structured representations in nonlinear models.
@@ 1 Introduction
Generative models have transformed information processing and demonstrated remarkable capacities for creativity in a variety of tasks ranging from vision to language to audio. The success of these models has been largely driven by modular, differentiable architectures based on deep neural networks that learn useful representations for downstream tasks. Recent years have seen increased interest in understanding and exploring the representations produced by these models through evolving lines of work on structured representation learning, identifiability and interpretability, disentanglement, and causal generative models. This work is motivated in part by the desire to produce performant generative models that also capture meaningful, semantic latent spaces that enable out-of-distribution generation under perturbations.
@@
A key conceptual driver of this line of work is the tradeoff between flexibility and structure, or expressivity and interpretability: It is widely believed that to learn structured, interpretable representations, model capacity must be constrained, sacrificing flexibility and expressivity. As described in Section 3, this intuition is supported by a growing body of theoretical work on nonlinear ICA (Hyvärinen and Pajunen, 1999), disentanglement (Bengio, 2013), and causal representation learning (Schölkopf et al., 2021). On the practical side, methods that have been developed to learn structured latent spaces tend to be bespoke to specific data types and models, and typically impose significant limitations on the flexibility and expressivity of the underlying models. Moreover, the most successful methods for learning structured representations typically impose fixed, known structure a priori, as opposed to learning this structure from data.
At the same time, there is a growing body of work that suggests generative models already learn surprisingly structured latent spaces (Mikolov et al., 2013; Szegedy et al., 2013; Radford et al., 2015, see also Section 3). This in turn suggests that existing models are already “close” to capturing the desired structure, and perhaps only small modifications are needed. Since we already have performant models that achieve state-of-the-art results in generation and discrimination, do we need to re-invent the wheel to achieve these goals? Our hypothesis is that we should be able to leverage the expressiveness of these models to build new models that learn latent structure from scratch. We emphasize that our goal is not to explain or interpolate the latent space of a pre-trained model, but rather to leverage known architectures to train a new model end-to-end from scratch.
@@
Figure 1: Overview of our approach. Given a black-box encoder-decoder architecture (blue), we propose to append a context module (red) to the head of the decoder. (left) Instead of passing the output of the encoder directly to the decoder (blue box + blue arrow), the embeddings $\mathbf{e}$ are passed through the context module, consisting of three distinct layers. The output of this module is then passed into the decoder. (right) The model learns to compose different concepts OOD, in this case object and background colour, to values that never appear in the training data.
@@
In this paper, we adopt this perspective: We start with a black-box encoder-decoder architecture, upon which we make no assumptions, and then augment this model to learn structured representations. The idea is that the black-box is an architecture that is already known to perform well on downstream tasks such as sampling and prediction, and so is flexible enough to capture complex patterns in data. We then introduce a modular, end-to-end differentiable architecture for learning structured representations, augmenting the original model with a simple decoder-only module that retains its existing capabilities while enabling concept identification and intervention as well as composing together multiple concept interventions for out-of-distribution (OOD) generation. The resulting latent structure is learned fully end-to-end, with no fixed structure imposed a priori. The motivation is to provide a framework for taking existing performant architectures and to train a new model that performs just as well, but with the added benefit of learning structured representations without imposing specific prior knowledge or structure a priori. These representations may be useful for interpretability, downstream tasks, or other purposes. Since the model, unlike previous approaches, imposes no latent structure a priori, we theoretically validate the approach with a novel identifiability result for structured representations recovered by the model.
To explore the implications of this modification, we run experiments including OOD generation. An attractive feature of our approach is that it makes running carefully controlled experiments and ablations particularly simple: We can carefully control hyperparameters across different models, ensuring that any differences in performance are attributable to specifically controlled architectural modifications. As a result, we contribute to a growing body of empirical work on the OOD capabilities of generative models by introducing a new set of ablations that can be used to guide the development of new methods.
@@
Our main contributions can be summarized as follows:
@@
    We introduce a simple module that can be attached to the head of an existing decoder block to learn concept representations by splitting each concept into a tensor slice, where each slice represents an interventional context in a reduced form structural equation model (SEM). This module can be attached to the head of any decoder block and trained end-to-end.
@@
    We conduct experiments on disentanglement and OOD generation on several benchmark datasets (3DIdent, CelebA, MNIST) as well as controlled simulations. To this end, we introduce a simple simulated visual environment for testing disentanglement and OOD generation. This allows us to carefully test different ablations under controlled settings.
@@
    To illustrate the adaptivity of our module to different architectures, we include experiments with several different VAE architectures including NVAE (Vahdat and Kautz, 2020).
As a matter of independent interest, we prove an identifiability result under concept interventions, and discuss how the proposed architecture can be interpreted as an approximation to this model. Full technical proofs are deferred to the appendix.
@@ 2 Preliminaries
Let $\mathbf{x}$ denote the observations, e.g., pixels in an image, and $\mathbf{z}$ denote hidden variables, e.g. latent variables that are to be inferred from the pixels. We are interested in training generative models with an encoder-decoder architecture, such as a variational autoencoder (VAE). A typical generative model consists of a decoder $p_{\theta}(\mathbf{x}\,|\,\mathbf{z})$ and an encoder $q_{\phi}(\mathbf{z}\,|\,\mathbf{x})$. After specifying a prior $p_{\theta}(\mathbf{z})$, this defines a likelihood by
[…]
@@
Both the decoder and encoder are specified by deep neural networks that will be trained end-to-end via standard techniques (Kingma and Welling, 2014; Rezende et al., 2014; Rezende and Mohamed, 2015). These networks are often chosen to capture implicit assumptions about the structure of the latent space, such as hierarchical or graphical structure, equivariance, etc. In this work, we focus on VAEs since they are often use to represent structured and semantic latent spaces, which is our focus. Unlike traditional structured VAEs, we do not impose a specific structure a priori.
Our starting point is the following incongruence: On the one hand, it is well-established that the latent spaces of VAEs are typically entangled and semantically misaligned, fail to generalize OOD, and suffer from posterior collapse (which has been tied to latent variable nonidentifiability, Wang et al., 2021). On the other hand, VAEs still learn highly structured latent spaces that can be traversed and interpolated, are “nearly” identifiable (Willetts and Paige, 2021; Reizinger et al., 2022), and represent abstract concepts linearly. Thus, our hypothesis is that sufficiently flexible models do capture meaningful structure internally, just not in a way that is interpretable or meaningful in practice. So, to learn latent structure, we build off the already performant embeddings of a generative model.
@@
Our set-up is the following: We are given a black-box VAE, consisting of an encoder and a decoder, on which we make no assumptions other than the encoder outputs a latent code corresponding to $\mathbf{x}$. In the notation of (2), this corresponds to $\mathbf{z}$, however, to distinguish the black-box embeddings from our model, we will denote the black-box embeddings hereafter by $\mathbf{e}$. Our plan is to work solely with the embeddings $\mathbf{e}$ and learn how to extract linear concept representations from $\mathbf{e}$ such that distinct concepts can be intervened upon and composed together to create novel, OOD samples. Crucially, we do not modify the black-box architecture in any way.
3 Related work
Structured generative models
To provide structure such as hierarchical, graphical, causal, and disentangled structures as well as other inductive biases in the latent space, there has been a trend towards building structured generative models that directly impose this structure a priori. Early work looked at incorporating fixed, known structure into generative models, such as autoregressive, graphical, and hierarchical structure (Germain et al., 2015; Johnson et al., 2016; Sønderby et al., 2016; Webb et al., 2018; Weilbach et al., 2020; Ding et al., 2021; Mouton and Kroon, 2023). This was later translated into known causal structure (Kocaoglu et al., 2017; Markham et al., 2023). When the latent structure is unknown, several techniques have been developed to learn useful (not necessarily causal) structure from data (Li et al., 2018; He et al., 2019; Wehenkel and Louppe, 2021; Kivva et al., 2022; Moran et al., 2022). More recently, based on growing interest in disentangled (Bengio, 2013) and/or causal (Schölkopf et al., 2021) representation learning, methods that learn causal structure have been developed (Markham and Grosse-Wentrup, 2020; Moraffah et al., 2020; Yang et al., 2021; Ashman et al., 2022; Shen et al., 2022; Kaltenpoth and Vreeken, 2023).
Causal representation learning
Causal representation learning (Schölkopf et al., 2021), which involves learning causal structure in the latent space, is a rapidly developing area that has produced fundamental results on theoretical aspects of identifiability (Brehmer et al., 2022; Shen et al., 2022; Lachapelle et al., 2022; Moran et al., 2022; Kivva et al., 2022; Buchholz et al., 2023; Gresele et al., 2021; Ahuja et al., 2023; von Kügelgen et al., 2023; Morioka and Hyvärinen, 2023; Liu et al., 2023; Mameche et al., 2023; Yao et al., 2023; Varici et al., 2023; Sturma et al., 2023; Xu et al., 2024; Zhang et al., 2024; Li et al., 2024; Varici et al., 2024; Bing et al., 2024; Ahuja et al., 2024; Talon et al., 2024; Yao et al., 2024). Recent work has also pushed in the direction of identifying concepts (Leemann et al., 2023; Rajendran et al., 2024; Fokkema et al., 2025). Our work draws inspiration from this line of work, which articulates precise conditions under which latent structure can be recovered in principle. By contrast, our focus is on methodological aspects of exploiting learned causal invariances in practice.
Linear representations
Generative models are known to represent concepts linearly in embedding space (e.g. Mikolov et al., 2013; Szegedy et al., 2013; Radford et al., 2015); see also Levy and Goldberg (2014); Arora et al. (2016); Gittens et al. (2017); Allen and Hospedales (2019); Ethayarajh et al. (2018); Seonwoo et al. (2019). This phenomenon has been well-documented over the past decade in both language models (Mikolov et al., 2013; Pennington et al., 2014; Arora et al., 2016; Conneau et al., 2018; Tenney et al., 2019; Elhage et al., 2022; Burns et al., 2022; Tigges et al., 2023; Nanda et al., 2023; Moschella et al., 2022; Li et al., 2023; Park et al., 2023; Gurnee et al., 2023; Jiang et al., 2024) and computer vision (Radford et al., 2015; Raghu et al., 2017; Bau et al., 2017; Engel et al., 2017; Kim et al., 2018; Chen et al., 2020; Wang et al., 2023; Trager et al., 2023). Our approach actively exploits this tendency by searching for concept representations as linear projections of the embeddings learned by a black-box model.
OOD generalization
A growing line of work studies the OOD generalization capabilities of generative models, with the general observation being that existing methods struggle to generalize OOD (Xu et al., 2022; Mo et al., 2024; Montero et al., 2022, 2021; Schott et al., 2022). It is worth noting that most if not all of this work evaluates OOD generalization using reconstruction on held-out OOD samples, as opposed to generation. For example, a traditional VAE may be able to reconstruct held-out samples, but it is not possible to actively sample OOD. See Section 5.1 for more discussion.
@@ 4 Architecture
We start with a black-box encoder-decoder architecture and assume there are $d_{c}$ concepts of interest, denoted by $\mathbf{c}=(\mathbf{c}_{1},\ldots,\mathbf{c}_{d_{c}})$. Our objectives are two-fold:
@@
    To learn structure between concepts from black-box embeddings $\mathbf{e}$ as linear projections $C\mathbf{e}$;
@@
    To compose concepts together in a single, transparent model that captures how different concepts are related.
We seek to learn these concepts in an end-to-end, differentiable manner.
A key intuition is that composition can be interpreted as a type of intervention in the latent space. This is sensible since interventions in a causal model are a type of nontrivial distribution shift. Thus, the problem takes on a causal flavour which we exploit to build our architecture. The difficulty with this from the causal modeling perspective is that encoding structural assignments and/or causal mechanisms directly into a feed-forward neural network is tricky, because edges between nodes within the same layer aren’t allowed in a feed-forward network but are required for the usual DAG representation of a causal model. To bypass this, we use the reduced form of a causal model which has a clear representation as a bipartite directed graph, with arrows only from exogenous to endogenous variables, making it conducive to being embedded into a feedforward NN. The tradeoff is that we do not learn a causal structure (i.e., a causal DAG), however, this enables the model to perform interventions directly in the latent space and to exploit invariances between interventional contexts.
@@ 4.1 Overview
[…]
Consistent with existing empirical work (Mikolov et al., 2013; Szegedy et al., 2013; Radford et al., 2015), we represent concepts as linear projections of the embeddings $\mathbf{e}$: Each concept $\mathbf{c}_{j}$ can be approximated as $\mathbf{c}_{j}\approx C_{j}\mathbf{e}$. This is modeled via a linear layer $\mathbf{c}\to\mathbf{e}$ that implicitly inverts this relationship in the decoder.
As a result, it makes sense to model the relationships between concepts with a linear SEM:
$\mathbf{c}_{j}=\sum_{k=1}^{d_{c}}\alpha_{kj}\mathbf{c}_{k}+\boldsymbol{\varepsilon}_{k},\quad\alpha_{kj}\in\mathbb{R}.$
By reducing this SEM, we deduce that
@@
We model this with a linear layer $\boldsymbol{\varepsilon}\to\mathbf{c}$, where the weights in this layer correspond to the matrix $A_{0}$. This layer will be used to encode the SEM between the concepts, which will be used to implement causal interventions.
@@
The final decoder architecture can be decomposed, at a high-level, as follows:

### v2

@@ Abstract
In designing generative models, it is commonly believed that in order to learn useful latent structure, we face a fundamental tension between expressivity and structure. In this paper we challenge this view by proposing a new approach to training arbitrarily expressive generative models that simultaneously learn causally disentangled concepts. This is accomplished by adding a simple context module to an arbitrarily complex black-box model, which learns to process concept information by implicitly inverting linear representations from the model’s encoder. Inspired by the notion of intervention in a causal model, our module selectively modifies its architecture during training, allowing it to learn a compact joint model over different contexts. We show how adding this module leads to causally disentangled representations that can be composed for out-of-distribution generation on both real and simulated data. The resulting models can be trained end-to-end or fine-tuned from pre-trained models. To further validate our proposed approach, we prove a new identifiability result that extends existing work on identifying structured representations.
keywords:
causal disentanglement; generative models; compositional generalization; concept learning; interventions.
@@ 1 Introduction
Generative models have transformed information processing and demonstrated remarkable capacities for creativity in a variety of tasks ranging from vision to language to audio. The success of these models has been largely driven by modular, differentiable architectures based on deep neural networks that learn useful representations for downstream tasks. Recent years have seen increased interest in understanding and exploring the representations produced by these models through evolving lines of work on structured representation learning, identifiability and interpretability, disentanglement, and causal generative models. This work is motivated in part by the desire to produce performant generative models that also capture meaningful, semantic latent spaces that enable out-of-distribution (OOD) generation.
@@
A key driver behind this work is the tradeoff between flexibility and structure, or expressivity and interpretability: Conventional wisdom suggests that to learn structured, interpretable representations, model capacity must be constrained, sacrificing flexibility and expressivity. This intuition is supported by a growing body of work on nonlinear ICA (Hyvärinen and Pajunen, 1999), disentanglement (Bengio, 2013), and causal representation learning (Schölkopf et al., 2021). On the practical side, methods that have been developed to learn structured latent spaces tend to be bespoke to specific data types and models, and typically impose limitations on the flexibility and expressivity of the underlying models. Moreover, the most successful methods for learning structured representations typically impose fixed, known structure a priori, as opposed to learning this structure from data.
@@
Figure 1: Overview (see (4) for notation map). Given a black-box encoder-decoder architecture (blue), we propose to prepend a context module (red) to the decoder. (left) Instead of passing the output of the encoder directly to the decoder (blue box + blue arrow), the embeddings $\mathbf{e}$ are passed through the context module, consisting of three distinct layers. The output of this module is then passed into the decoder. (right) The model learns to compose different concepts OOD, e.g., object and background colour, to values that never appear together in the training data.
At the same time, there is also a growing body of work that suggests generative models already learn surprisingly structured latent spaces (e.g. Mikolov et al., 2013; Szegedy et al., 2013; Radford et al., 2015). This in turn suggests that existing models are already “close” to capturing useful structure, and perhaps only small modifications are needed. Since we already have performant models that achieve state-of-the-art results in generation and prediction, do we need to re-invent the wheel to achieve these goals? Our hypothesis is that we should be able to leverage the expressiveness of these models to build new models that learn latent structure from scratch. We emphasize that our goal is not to explain or interpolate the latent space of a pre-trained model, but rather to leverage known architectures to train a new model, either end-to-end or by fine-tuning an existing model.
@@
In this paper, we adopt this perspective: We start with an arbitrarily complex black-box model, upon which we make no assumptions, then augment this model to learn causally disentangled representations. The idea is that the black-box model is already known to perform well on downstream tasks (e.g., generation, prediction), and so is flexible enough to capture complex patterns in data. We introduce a modular, end-to-end differentiable architecture for learning causally disentangled representations (Bengio, 2013; Schölkopf et al., 2021) that augments the original model with a simple module that retains its existing capabilities while enabling concept identification and intervention as well as composing together multiple concept interventions for OOD generation (Figure 1). The resulting latent structure is learned entirely from data, with no fixed structure imposed a priori. The motivation is to provide a framework for taking existing performant architectures and to train a new model that performs just as well, but with the added benefit of learning structured representations without imposing specific prior knowledge or structure. Since the model, unlike previous approaches, does not impose rigid latent structure, we theoretically validate the approach with a novel identifiability result for structured representations recovered by the model. For a detailed discussion of related work, see Appendix A.
@@
More specifically, we make the following contributions:
@@
    We introduce a context module that is attached to a decoder to learn concept representations by splitting each concept into a tensor slice, where each slice represents an interventional context in a reduced form structural equation model (SEM). This module can be attached to any decoder and trained end-to-end or fine-tuned (Figure 1).
@@
    We evaluate this module through quantitative and qualitative experiments on OOD generation, which is distinct from and significantly more challenging than OOD reconstruction as in prior work (e.g. Xu et al., 2022; Mo et al., 2024; Montero et al., 2022; Schott et al., 2022). We perform experiments on OOD generation on several real datasets as well as carefully controlled simulations.
@@
    We illustrate the adaptivity of our module to different architectures (including NVAE, Vahdat and Kautz, 2020) and against multiple ablations that control model capacity, data contexts, training protocols, and hyperparameters, ensuring that any differences in performance are directly attributable to the context module.
  4. 4.
    We fine-tune existing, pre-trained models after attaching the context module, giving substantial computational savings compared to end-to-end training while improving the learned representations and lending OOD generation abilities to pre-trained models.
As a matter of independent interest, we prove an identifiability result under concept interventions (Theorem 3.3), and discuss how the proposed architecture can be interpreted as an approximation to this model. We also introduce quad, a simple simulated visual environment for evaluating causal disentanglement, compositional abilities, and OOD generation. Due to space limitations, full technical proofs, along with detailed literature review and experimental details, can be found in the appendix.
@@ 2 Preliminaries
Let $\mathbf{x}$ denote the observations, e.g., pixels in an image, and $\mathbf{z}$ denote hidden variables, e.g., latent variables that are to be inferred from the pixels. A typical generative model consists of a decoder $p_{\theta}(\mathbf{x}\,|\,\mathbf{z})$ and an encoder $q_{\phi}(\mathbf{z}\,|\,\mathbf{x})$. After specifying a prior $p_{\theta}(\mathbf{z})$, this defines a likelihood by
$p_{\theta}(\mathbf{x})=\int p_{\theta}(\mathbf{x}\,|\,\mathbf{z})p_{\theta}(\mathbf{z})\operatorname{d\!}{\mathbf{z}}.$ (1)
@@
Both the decoder and encoder are specified by deep neural networks that are trained end-to-end via standard techniques (Kingma and Welling, 2014; Rezende et al., 2014; Rezende and Mohamed, 2015). In this work, we focus on variational autoencoders (VAEs) since they are often use to represent structured and semantic latent spaces, which is our focus. Unlike traditional structured VAEs, we do not impose a specific structure a priori.
Causal disentanglement
Our goal is to learn distinct latent factors that can be composed together. This is closely related to causal disentanglement, which we recall informally here:
Definition 2.1 (Bengio, 2013; Thomas et al., 2017).
A causally disentangled latent space consists of separately controllable (intervenable) latent factors.
In particular, we want to manipulate and intervene on latent factors in a causally meaningful way; crucially this allows us to compose distinct concepts together without needing to learn the full causal graph in the latent space. This notion of disentanglement is also different from the axis-alignment notion (Kim and Mnih, 2018; Chen et al., 2018). The “causal” aspect comes from interpreting “separately controllable” to mean “able to intervene on independent causal mechanisms”, as suggested by Locatello et al. (2019); see also Schölkopf et al. (2021). See Appendix B for formal details. Rather than attempt to reconstruct the full causal DAG model (as in Yang et al., 2021; Squires et al., 2023; Buchholz et al., 2023), we aim for the more practical goal of learning different interventional contexts through distributional invariances; thus our approach strikes a balance between easy-to-learn unstructured latent spaces and hard-to-learn but desirable fully structured causal DAG latent spaces.
Measuring causal disentanglement
A standard approach to quantifying causal disentanglement is to evaluate the recovery of the latent causal graph; since our explicit goal is to avoid learning this graph, we propose a different metric that is more suitable to downstream applications. To measure causal disentanglement, we use OOD generation on novel interventions as a metric. We train models so that they never see examples of certain latent factors composed together (equivalently, multiple simultaneous interventions on distinct factors), so these examples are genuinely OOD. By withholding these examples during training and building a model that is capable of intervening on multiple factors, we can compare the quality of OOD samples generated from our model to the held-out OOD samples. In our experiments, we use the sliced Wasserstein (SW) metric (18); see Appendix D.1 for details.
OOD reconstruction vs. OOD generation
Previous work has evaluated the OOD capabilities of generative models by using OOD reconstruction as a metric (e.g. Xu et al., 2022; Mo et al., 2024; Montero et al., 2022; Schott et al., 2022). There is a crucial difference between OOD reconstruction and OOD generation: By OOD generation, we mean the ability to conditionally generate and combine novel interventions on existing concepts. For example, suppose during training the model only sees small, red objects along with big, blue objects. At test time, we wish to generate OOD samples of big, red objects or small, blue objects that were never seen during training.
The crucial difference is that while we can always attempt to reconstruct any sample (i.e., OOD or not) and evaluate its reconstruction error, it is not always possible to have the model generate a new OOD sample. Unless the model learns specific latent factors corresponding to concepts such as size or colour, we cannot control (i.e., intervene) the size or colour of random samples from the model.
Thus, OOD generation not only evaluates the ability of a model to compose learned concepts in new ways, but also the ability of a model to identify and capture underlying concepts of interest. Accordingly, we argue that OOD generation is more appropriate to evaluate causal disentanglement since it captures both the model’s ability to learn concepts as distinct latent factors as well as compose them together in novel ways.
Set-up and approach
@@
Our set-up is the following: We have a black-box encoder and decoder, on which we make no assumptions other than the encoder outputs a latent code corresponding to $\mathbf{x}$. In the notation of (1), this corresponds to $\mathbf{z}$, however, to distinguish the black-box embeddings from our model, we will denote the black-box embeddings hereafter by $\mathbf{e}$. Our plan is to work solely with the embeddings $\mathbf{e}$ and learn how to extract linear concept representations from $\mathbf{e}$ such that distinct concepts can be intervened upon and composed together to create novel, OOD samples.
Our approach is motivated by the following incongruence: On the one hand, it is well-established that the latent spaces of generative models are typically entangled and semantically misaligned, fail to generalize OOD, and suffer from posterior collapse (which has been tied to latent variable nonidentifiability, Wang et al., 2021). On the other hand, they still learn structured latent spaces that can be interpolated, are “nearly” identifiable (Willetts and Paige, 2021; Reizinger et al., 2022), and represent abstract concepts linearly (Mikolov et al., 2013; Szegedy et al., 2013; Radford et al., 2015). So, to learn latent structure, we simply need to extract the right concept representations from the already performant embeddings of existing architectures and learn how to compose them together.
@@ 3 Architecture
We start with a black-box encoder-decoder architecture along with $d_{c}$ concepts of interest, denoted by $\mathbf{c}=(\mathbf{c}_{1},\ldots,\mathbf{c}_{d_{c}})$. Our objectives are three-fold:
@@
    To extract concepts from black-box embeddings $\mathbf{e}$ as linear projections $C\mathbf{e}$;
@@
    To compose concepts together in a single, transparent model that captures how different concepts are related;
  3. 3.
    To accomplish both of these objectives without restricting the black-box encoder or decoder architectures, and without sacrificing sampling quality.
A key intuition is that composition can be interpreted as a type of intervention in the latent space. This is sensible since interventions in a causal model are a type of nontrivial distribution shift. Thus, the problem takes on a causal flavour which we leverage to build our architecture. The difficulty with this from the causal modeling perspective is that encoding structural assignments and/or causal mechanisms directly into a feed-forward neural network is tricky, because edges between nodes within the same layer are not allowed in a feed-forward network but are required for the usual DAG representation of a causal model. To bypass this, we use the reduced form of a causal model which has a clear representation as a bipartite directed graph, with arrows only from exogenous to endogenous variables, making it conducive to being embedded into a neural network. The tradeoff is that this does not learn a causal DAG, however it still enables the model to perform interventions directly in the latent space and to leverage invariances between interventional contexts.
@@ 3.1 Overview
$\mathbf{x}\overset{\textup{encoder}}{\xrightarrow{\hskip 16.38895pt}}\mathbf{e}\overset{\textup{decoder}}{\xrightarrow{\hskip 16.38895pt}}\widehat{\mathbf{x}}$.
Based on an extensive body of empirical work that shows concepts are linearly represented (Mikolov et al., 2013; Szegedy et al., 2013; Radford et al., 2015, see Appendix A for more discussion), we represent concepts as linear projections of the embeddings $\mathbf{e}$: Each concept $\mathbf{c}_{j}$ can be approximated as $\mathbf{c}_{j}\approx C_{j}\mathbf{e}$. This is modeled via a linear layer $\mathbf{c}\to\mathbf{e}$ that implicitly inverts this relationship in the decoder.
We further model the relationships between concepts with a linear SEM:
$\mathbf{c}_{j}=\sum_{k=1}^{d_{c}}\alpha_{kj}\mathbf{c}_{k}+\boldsymbol{\varepsilon}_{k},\quad\alpha_{kj}\in\mathbb{R}.$ (2)
By reducing this SEM and solving for $\mathbf{c}$, we deduce that
@@
This is known as the reduced form of the SEM (2). We model this with a linear layer $\boldsymbol{\varepsilon}\to\mathbf{c}$, where the weights in this layer correspond to the matrix $A_{0}$. This layer will be used to encode the SEM between the concepts, which will be used to implement causal interventions.
@@
The final decoder architecture can be decomposed at a high-level as follows (Figure 1):